# Libra: Leveraging Temporal Images for Biomedical Radiology Analysis

### v1

@@ B.3 Evaluation Metrics Details
RadGraph-based metrics The RadGraph model (Jain et al., 2021) is designed to parse radiology reports into structured graphs. These graphs consist of clinical entities, which include references to anatomy and observations, as well as the relationships between these entities. This structured representation enables a more detailed and systematic analysis of radiology reports, facilitating downstream tasks such as information extraction, report generation, and clinical decision support.
@@
CheXpert F1 This set of metrics utilizes the CheXbert automatic labeler (Smit et al., 2020) to extract “present”, “absent”, or “uncertain” labels for each of the 14 CheXpert pathologies (Irvin et al., 2019a) from the generated reports and their corresponding references. In line with prior work, we report CheXpert-F1 for all 14 classes, as well as for the 5 most common findings in real-world CXR reports, referring to these as ‘[Macro/Micro]-F1-[5/14]’.
CheXbert vector similarity We also employ CheXbert vector similarity (Yu et al., 2022), which calculates the cosine similarity between the embeddings of the generated and reference reports after processing them through the CheXbert model (Smit et al., 2020).
RadCliQ In addition, we utilise RadCliQ (Radiology Report Clinical Quality), a composite metric that combines RadGraph-F1 and BLEU scores in a linear regression model to estimate the number of errors that radiologists are likely to detect in a report. To maintain consistency with previous research, we use version of it.
@@
Temporal Entity F1 The , which is specifically designed to detect temporal entities. This score is neither a traditional lexical metric nor a adiology-specific metrics. Here is an example to help you understand better, as shown in Table 6.
@@
From these results, it is clear that the differences in lexical (ROUGE-L) and clinical (RadGraph-F1) metrics between the two candidates are relatively smaller compared to the score. This highlights that Temporal Entity F1 effectively captures and evaluates the quality of temporal information in radiology reports, distinguishing it more accurately than other standard metrics in the context of temporal information descriptions.
B.4 Prompt Example
Here, we select examples from the MIMIC-CXR (Johnson et al., 2019b) dataset as instances and further synthesise them using GPT-4 (OpenAI et al., 2024) to avoid any ethical concerns, as shown in Table 7. Following Hyland et al. (2024)’s work, we adopt a rule-based method to extract relevant sections from the report of the current image, utilising a fixed system prompt and a dynamic clinical prompt for each current scan. We utilised four clinical instructions from the original report: {Indication}, {History}, {Comparison}, and {Technique}. Notably, unlike MAIRA-2 (Bannur et al., 2024), we do not utilise the report of the prior image.
| Original Radiology Report |
| EXAMINATION: Chest (Portable AP) |
| INDICATION: Dyspnea and cough, right-sided back pain. |
| HISTORY: Intubation with pulmonary edema. |
| TECHNIQUE: Portable upright chest radiograph. |
| Prompt Content |
| [System prompt] { |
| Image Representation Placeholder |
| [Clinical prompt] { |
| Indication: Dyspnea and cough, right-sided back pain. |
| History: Intubation with pulmonary edema. |
| Technique: Portable upright chest radiograph. |
Appendix C Appendix C
@@
C.1 Impact of the Temporal Alignment Connector under General-Domain Pre-trained Models
@@
Domain-specific pre-trained models (i.e., RAD-DINO (Pérez-García et al., 2024) and Meditron (Chen et al., 2023b)) inherently incorporate domain-specific knowledge, such as phrasing conventions, pronoun usage, and even temporal information embedded in the training corpus.
To eliminate the influence of such factors, we used a general-domain image encoder (DINOv2 (Oquab et al., 2024)) and a LLM (Vicuna-7B-v1.5 (Chiang et al., 2023)), allowing the structural enhancements of TAC to be observed more directly. We adopted the same setup as the second ablation study in Section 3. We first conducted a baseline experiment, referred to as Libra-b, by fine-tuning only the adapter for the Findings generation task. As shown in Table 8, we then conducted ablation studies by sequentially removing different components from the model, including the TFM, LFE, Prior Image Prefix Bias (PIPB) in TFM, and the entire TAC. Removing TFM restricts the model to processing only the current image, using a configuration similar to LLaVA (Liu et al., 2023), but with a four-layer MLP to align the image representation with the LLM’s hidden dimensions. It is important to note that without the TFM module, the model cannot process true prior images or dummy prior images, and is limited to only the current image as input. Without LFE, the model follows the LLaVA setup, using the penultimate layer of the image encoder to process single or paired images.
The ablation results are consistent with those observed when using domain-specific pre-trained models, as presented in Section 3. Compared to Libra-b, removing any TAC submodule results in a decline across all metrics. Removing the TFM caused a notable drop in the F1 score (4%), emphasising its role in capturing temporal information. The removal of the LFE significantly decreased RadGraph-related scores, highlighting its importance for extracting detailed image features. Removing the PIPB impacted clinical scores more than lexical scores. Finally, removing the TAC resulted in significant declines across all metrics, underscoring its critical role in integrating image details and temporal information.
C.2 Impact of the Temporal Alignment Connector After the Second Stage
@@
To further evaluate the impact of the Temporal Alignment Connector (TAC) on Libra’s performance, we followed the setup of the second ablation study in Section 3. After the first stage of alignment, the model underwent a second stage of fine-tuning. This stage was designed to optimise the model’s performance on the Findings section generation task by leveraging the aligned visual and textual features learned during the initial stage.
In this phase, we applied Low-Rank Adaptation (LoRA) (Hu et al., 2021) to fine-tune the pre-trained LLM (Meditron (Chen et al., 2023b)), while keeping the visual encoder (RAD-DINO (Pérez-García et al., 2024)) and TAC weights frozen. The baseline for this experiment is Libra-2 (in Table 9), which is derived from Libra-1 (in Table 4) after undergoing LoRA fine-tuning.
We conducted ablation studies by progressively removing different TAC components, including TFM, LFE, the Prior Image Prefix Bias (PIPB), and the entire TAC. The conclusions remain consistent with those in Section 3, demonstrating that removing any TAC submodule results in a decline across all metrics compared to Libra-2. This indicates that the performance improvements brought by TAC are stable and unaffected by changes in training stages. It further confirms that TAC has effectively embedded the capability to process temporal information within the model.
C.3 Robustness Assessment of the Temporal Alignment Connector
@@
In this experiment, we introduced an additional round of LoRA fine-tuning to induce overtraining and assess the robustness of the Temporal Alignment Connector (TAC). Following the setup in Appx. C.2, after integrating the first LoRA weights, we reinitialised a new set of LoRA adapters for the LLM and conducted further training under the same second-stage fine-tuning configuration for one epoch. The baseline for this experiment is Libra-3 (as shown in Table 10), which is derived from Libra-2 (illustrated in Table 9) following this additional fine-tuning step.
The results reveal that, compared to Libra-2, Libra-3 demonstrates minimal changes in lexical scores, while clinical scores decline due to overfitting caused by the additional fine-tuning. Notably, the CheXpert (Smit et al., 2020) (Macro-F1-[5/14]) scores exhibit the most influential reduction.
However, the ablation studies reaffirm that the performance improvements introduced by TAC are robust and remain unaffected by variations in training strategies. This robustness is attributed to TAC’s ability to effectively capture and preserve temporal image representations during the first training phase, which are subsequently preserved throughout further fine-tuning.
These findings highlight that TAC reliably embeds the capacity for temporal information processing into the model, making it a stable and essential component for handling temporal data in RRG tasks.
C.4 Impact of Radiology-Specific Pre-trained Models

### v2

@@ F.2 Clinical Metrics
RadGraph-based metrics
RadGraph model (Jain et al., 2021) is designed to parse radiology reports into structured graphs. These graphs consist of clinical entities, which include references to anatomy and observations, as well as the relationships between these entities. This structured representation enables a more detailed and systematic analysis of radiology reports, facilitating downstream tasks such as information extraction, report generation, and clinical decision support.
@@
CheXpert F1
This set of metrics utilizes the CheXbert automatic labeler (Smit et al., 2020a) to extract “present”, “absent”, or “uncertain” labels for each of the 14 CheXpert pathologies (Irvin et al., 2019) from the generated reports and their corresponding references. In line with prior work, we report CheXpert-F1 for all 14 classes, as well as for the 5 most common findings in real-world CXR reports, referring to these as “[Macro/Micro]-F1-[5/14]”.
CheXbert vector similarity
RadCliQ
In addition, we utilise RadCliQ (Radiology Report Clinical Quality) (Yu et al., 2022), a composite metric that combines RadGraph-F1 and BLEU scores in a linear regression model to estimate the number of errors that radiologists are likely to detect in a report. To maintain consistency with previous research, we use version of it.
@@
F.3 Temporal Entity F1
We introduced , a metric specifically designed to detect temporal entities reflecting changes over time. Unlike traditional lexical or radiology-specific metrics, evaluates the quality of temporal information in radiology reports.
@@ F.3 Temporal Entity F1
As shown in Table 6, the differences in lexical (ROUGE-L (Lin, 2004)) and clinical (RadGraph-F1 (Jain et al., 2021)) metrics between the two candidates are relatively smaller compared to the score. This demonstrates that Temporal Entity F1 effectively captures and evaluates the quality of temporal information in radiology reports, distinguishing it more accurately than other standard metrics in the context of temporal information descriptions.
Appendix G Analysis of Concurrent Work and Non-LLM-based Models
G.1 Discussion on Performance with Radiology Foundation Models
| Metric | RaDialog | MedVersa | MAIRA-2 | M4CXR | Libra (%) |
| Lexical: | |||||
| ROUGE-L | 31.6 | – | 38.4 | 28.5 | 36.7 (-4.4%) |
| BLEU-1 | 39.2 | – | 46.5 | 33.9 | 51.3 (10.3%) |
| BLEU-4 | 14.8 | 17.8 | 23.4 | 10.3 | 24.5 (4.7%) |
| METEOR | – | – | 42.0 | – | 48.9 (16.4%) |
| BERTScore | – | 49.7 | – | – | 62.5 (25.8%) |
| Clinical: | |||||
| RadGraph-F1 | – | 28.0 | 34.6 | 21.8 | 32.9 (-4.9%) |
| RG | – | – | 39.7 | – | 37.6 (-5.3%) |
| RadCliQ() | – | 2.7 | 2.6 | – | 2.7 (-3.8%) |
| CheXbert vector | – | 46.4 | 50.6 | – | 46.9 (-7.3%) |
| CheXpert-F1: | |||||
| Micro-F1-14 | 39.2 | – | 58.5 | 60.6 | 55.9 (-7.8%) |
| Macro-F1-14 | – | – | 42.7 | 40.0 | 40.4 (-5.4%) |
| Micro-F1-5 | – | – | 58.9 | 61.8 | 60.1 (-2.8%) |
| Macro-F1-5 | – | – | 51.5 | 49.5 | 53.8 (4.5%) |
As shown in Table 7, these models belong to the category of radiology foundation models.
DaDialog (Pellegrini et al., 2023) is a conversational MLLM designed for a broad range of dialogue-based medical assistance tasks. To enhance structured findings extraction, it employs the publicly available CheXbert model (Smit et al., 2020b) to extract symptom labels from scans, facilitating a structured representation of findings.
MedVersa (Zhou et al., 2024) and M4CXR (Park et al., 2024) support a diverse set of tasks, including medical report generation, visual grounding, and visual question answering. These models aim to provide general-purpose multimodal medical assistance by leveraging vision-language pre-training strategies.
MAIRA-2 (Bannur et al., 2024) specialises in grounded radiology report generation, which differs from traditional report generation tasks by requiring explicit image-level localization of findings and symptoms. Grounded radiology reporting, as defined by Bannur et al. (2024), structures the report as a list of sentences, where each sentence: (1) is linked to zero or more spatial image annotations, and (2) describes at most a single finding from an image. To support this task, MAIRA-2 introduces a custom dataset, explicitly designed to provide structured annotations aligning textual descriptions with spatial regions of interest in radiological images. This approach contrasts with conventional RRG models that generate unstructured free-text reports.
It is worth noting that the inference sets differ slightly across these models. Additionally, all these models leverage supplementary radiology information, such as lateral view scans, prior study reports, or both (as detailed in Appx. A.2), to enhance their performance in radiology-related tasks.
G.2 Discussion on Performance with non-LLM-based Models
| Model | Lexical Metrics | Clinical Metrics | |||||||
| B-1 | B-2 | B-3 | B-4 | MTR | R-L | P | R | F1 | |
| ST | 29.9 | 18.4 | 12.1 | 8.4 | 12.4 | 26.3 | 24.9 | 20.3 | 20.4 |
| ATT2IN | 32.5 | 20.3 | 13.6 | 9.6 | 13.4 | 27.6 | 32.3 | 23.9 | 20.4 |
| ADAATT | 29.9 | 18.5 | 12.4 | 8.8 | 11.8 | 26.6 | 26.8 | 18.6 | 18.1 |
| TopDown | 31.7 | 19.5 | 13.0 | 9.2 | 12.8 | 26.7 | 32.0 | 23.1 | 23.8 |
| R2Gen | 35.3 | 21.8 | 14.5 | 10.3 | 14.2 | 27.0 | 33.3 | 27.3 | 27.6 |
| R2GenCMN | 35.3 | 21.8 | 14.8 | 10.6 | 14.2 | 27.8 | 34.4 | 27.5 | 27.8 |
| XPRONET | 34.4 | 21.5 | 14.6 | 10.5 | 13.8 | 27.9 | – | – | – |
| CMCL | 34.4 | 21.7 | 14.0 | 9.7 | 13.3 | 28.1 | – | – | – |
| PPKED | 36.0 | 22.4 | 14.9 | 10.6 | 14.9 | 28.4 | – | – | – |
| AlignTransformer | 37.8 | 23.5 | 15.6 | 11.2 | 15.8 | 28.3 | – | – | – |
| CA | 35.0 | 21.9 | 15.2 | 10.9 | 15.1 | 28.3 | 35.2 | 29.8 | 30.3 |
| LKBMA | 38.6 | 23.7 | 15.7 | 11.1 | – | 27.4 | 42.0 | 33.9 | 35.2 |
| M2TR | 37.8 | 23.2 | 15.4 | 10.7 | 14.5 | 27.2 | 24.0 | 42.8 | 30.8 |
| KnowMAT | 36.3 | 22.8 | 15.6 | 11.5 | – | 28.4 | 45.8 | 34.8 | 37.1 |
| RAMT | 36.2 | 22.9 | 15.7 | 11.3 | 15.3 | 28.4 | 38.0 | 34.2 | 33.5 |
| CMM-RL | 38.1 | 23.2 | 15.5 | 10.9 | 15.1 | 28.7 | 34.2 | 29.4 | 29.2 |
| CMCA | 36.0 | 22.7 | 15.6 | 11.7 | 14.8 | 28.7 | 44.4 | 29.7 | 35.6 |
| KiUT | 39.3 | 24.3 | 15.9 | 11.3 | 16.0 | 28.5 | 37.1 | 31.8 | 32.1 |
| DCL | – | – | – | 10.9 | 15.0 | 28.4 | 47.1 | 35.2 | 37.3 |
| MMTN | 37.9 | 23.8 | 15.9 | 11.6 | 16.1 | 28.3 | – | – | – |
| METrans | 25.0 | 16.9 | 12.4 | 15.2 | – | 29.1 | 36.4 | 30.9 | 31.1 |
| ORGAN | 38.6 | 25.6 | 17.2 | 12.3 | 16.2 | 29.3 | 41.6 | 41.8 | 38.5 |
| COMG | 36.3 | 23.5 | 16.7 | 12.4 | 12.8 | 29.0 | – | – | – |
| MedM2G | 41.2 | 26.9 | 17.9 | 14.2 | – | 30.9 | – | – | – |
| CvT2DistilGPT2 | 39.2 | 24.5 | 16.9 | 12.4 | 15.3 | 28.5 | 35.9 | 41.2 | 38.4 |
| RGRG | 37.3 | 24.9 | 17.5 | 12.6 | 16.8 | 26.4 | 46.1 | 47.5 | 44.7 |
| BioViL-T | – | – | – | 9.2 | – | 29.6 | – | – | 17.5 |
| VLCI | 40.0 | 24.5 | 16.5 | 11.9 | 15.0 | 28.0 | 48.9 | 34.0 | 40.1 |
| TiBiX | 32.4 | 23.4 | 18.5 | 15.7 | 16.2 | 33.1 | 30.0 | 22.4 | 25.0 |
| RECAP | 42.9 | 26.7 | 17.7 | 12.5 | 16.8 | 28.8 | 38.9 | 44.3 | 39.3 |
| MS-TF | 43.6 | 27.5 | 18.4 | 12.9 | 17.7 | 30.5 | – | – | 41.1 |
| Libra | 51.3 | 38.0 | 30.0 | 24.5 | 48.9 | 36.7 | 59.7 | 52.5 | 55.9 |
To facilitate comparison with non-LLM-based models, we selected evaluation metrics commonly used in these studies. These include BLEU- (Papineni et al., 2002), METEOR (MTR) (Banerjee and Lavie, 2005), and ROUGE-L (R-L) (Lin, 2004). For clinical metrics, we adopted the CheXbert (Irvin et al., 2019), reporting Precision (P), Recall (R), and F1.
Baseline
For performance evaluation, we compare our model with the following baselines: ST (Vinyals et al., 2015), ATT2IN (Rennie et al., 2017), ADAATT (Lu et al., 2017), TopDown (Anderson et al., 2018), R2Gen (Chen et al., 2020), R2GenCMN (Chen et al., 2021), M2TR (Nooralahzadeh et al., 2021), CMCL (Liu et al., 2021a),PPKED (Liu et al., 2021b), AlignTransformer (You et al., 2021), CA (Liu et al., 2021c), LKBMA (Yang et al., 2022b), KnowMAT (Yang et al., 2022a), XPRONET (Wang et al., 2022), CMM-RL (Qin and Song, 2022), RAMT (Zhang et al., 2024b), CMCA (Song et al., 2022), KiUT (Huang et al., 2023), DCL (Li et al., 2023b), MMTN (Cao et al., 2023), METrans (Wang et al., 2023), ORGAN (Hou et al., 2023b), COMG (Gu et al., 2023), BioViL-T (Bannur et al., 2023), RGRG (Tanida et al., 2023), RECAP (Hou et al., 2023a), CvT2DistilGPT2 (Nicolson et al., 2023), VLCI (Chen et al., 2024a), TiBiX (Sanjeev et al., 2024), MedM2G (Zhan et al., 2024), MS-TF (Mei et al., 2024).
To ensure fairness, Libra also utilizes prior images, aligning with other models that leverage prior images or additional information. As demonstrated in Table 8, Libra, similar to other LLM-based models, consistently outperforms non-LLM-based models. This advantage is largely attributed to advancements in LLMs and visual instruction tuning (Liu et al., 2023), enabling multimodal large language models (MLLMs) to achieve superior performance in RRG tasks.
Appendix H Additional Ablation Studies
H.1 Impact of Temporal Information on Libra in RRG
| Metric | Libra | |
| w/o prior | w/ prior (%) | |
| Lexical: | ||
| ROUGE-L | 36.17 | 36.66 (+1.35%) |
| BLEU-1 | 51.20 | 51.25 (+0.10%) |
| BLEU-4 | 24.33 | 24.54 (+0.86%) |
| METEOR | 48.69 | 48.90 (+0.43%) |
| BERTScore | 61.94 | 62.50 (+0.90%) |
| F1 | 32.72 | 35.34 (+8.00%) |
| Clinical: | ||
| RadGraph-F1 | 32.42 | 32.87(+1.39%) |
| RG | 36.92 | 37.57(+1.76%) |
| RadCliQ() | 2.76 | 2.72 (+1.45%) |
| CheXbert vector | 46.31 | 46.85 (+1.17%) |
| CheXpert-F1: | ||
| Micro-F1-14 | 55.25 | 55.87 (+1.12%) |
| Macro-F1-14 | 40.15 | 40.38(+0.57%) |
| Micro-F1-5 | 58.93 | 60.07(+1.93%) |
| Macro-F1-5 | 52.61 | 53.75(+2.17%) |
Temporal information is embedded in paired images and referenced in the corresponding radiology reports, capturing changes over time through references to prior symptoms and their progression, as discussed in Appx. B.1. As shown in Table 5, 86% of the test data includes prior images, providing a solid foundation for evaluating the impact of temporal information.
During training, Libra integrates the ability to perceive and utilise temporal information into its architecture. To evaluate whether Libra effectively leverage temporal information during inference, we assess its performance using prior images when available as references to determine their impact on the overall capability.
In Table 9, the inclusion of prior images substantially enhances Libra’s performance across all metrics. Notably, clinical scores exhibit greater improvements compared to lexical scores, underscoring the importance of temporal information in generating high-quality medical reports beyond merely improving linguistic fluency.
The F1 score shows the most substantial improvement, with an increase of 8%, highlighting Libra’s capability to effectively leverage temporal changes provided by prior images. These results validate the role of temporal information in enhancing the quality of the generated Findings section and improving Libra’s overall performance in RRG tasks.
@@ H.1 Impact of Temporal Information on Libra in RRG
H.2 Impact of the Temporal Alignment Connector under General-Domain Pre-trained Models
@@
Domain-specific pre-trained models (i.e., RAD-DINO (Pérez-García et al., 2024) and Meditron (Chen et al., 2023c)) inherently incorporate domain-specific knowledge, such as phrasing conventions, pronoun usage, and even temporal information embedded in the training corpus. To isolate the structural impact of TAC, we used a general-domain image encoder (DINOv2 (Oquab et al., 2024)) and a LLM (Vicuna-7B-v1.5 (Chiang et al., 2023)), allowing the structural enhancements of TAC to be observed more directly.
We replicated the first ablation setup from Sec. 4. We first conducted a baseline experiment, referred to as Libra-b, by fine-tuning only the adapter for the Findings generation task. As shown in Table 10, we then conducted ablation studies by sequentially removing different components from the model, including the Temporal Fusion Module (TFM), Layerwise Feature Extractor (LFE), Prior Image Prefix Bias (PIPB), and the entire TAC. Removing TFM restricts the model to processing only the current image, using a configuration similar to LLaVA (Liu et al., 2023), but with a four-layer MLP to align the image feature with the LLM’s hidden dimensions. Notably, without TFM, the model cannot process prior images or dummy prior images, and is limited to only the current image as input. Without LFE, the model follows the LLaVA setup, using the penultimate layer of the image encoder to process single or paired images.
The ablation results are consistent with those observed using domain-specific models, as presented in Table 2. Removing any TAC submodule led to declines across all metrics. Specifically, removing TFM caused a notable drop in the F1 score (4%), emphasising its role in capturing temporal information. The absence of LFE significantly reduced RadGraph-related scores, demonstrating its importance for detailed image feature extraction. PIPB removal primarily impacted clinical metrics, while removing the entire TAC resulted in substantial declines across all metrics. These findings reaffirm the critical role of TAC in integrating image details and temporal information effectively.
H.3 Impact of the Temporal Alignment Connector After the Second-Stage Fine-tuning
@@
To further evaluate the impact of the Temporal Alignment Connector (TAC) on Libra’s performance, we followed the setup of the first ablation study in Sec. 4. After the first stage of alignment, the model underwent a second stage of fine-tuning. This stage was designed to optimise the model’s performance on the Findings section generation task by leveraging the aligned visual and textual features learned during the initial stage.
In this phase, we applied Low-Rank Adaptation (LoRA) Hu et al. (2021) to fine-tune the pre-trained LLM (Meditron Chen et al. (2023c)), while keeping the visual encoder (RAD-DINO Pérez-García et al. (2024)) and TAC weights frozen. The baseline for this experiment is Libra-2 (in Table 11), which is derived from Libra-1 (in Table 2) after undergoing LoRA fine-tuning.
We conducted ablation studies by progressively removing different TAC components, including TFM, LFE, the Prior Image Prefix Bias (PIPB), and the entire TAC. Results consistently showed declines across all metrics compared to Libra-2, mirroring the trends observed in Sec. 4. This reinforces that the performance improvements brought by TAC are stable and unaffected by changes in training stages. It further confirms that TAC has embedded the capability to process temporal information within the model.
H.4 Robustness Evaluation of the Temporal Alignment Connector
To evaluate the robustness of the Temporal Alignment Connector (TAC), we introduced an additional round of LoRA fine-tuning to induce overtraining. Following the setup in Appx. H.3, after integrating the first LoRA weights, a new set of LoRA adapters was reinitialised for the LLM and trained for one epoch under the same second-stage fine-tuning configuration. The baseline for this experiment is Libra-3 (as shown in Table 12), which is derived from Libra-2 (illustrated in Table 11) following this additional fine-tuning step.
@@ H.4 Robustness Evaluation of the Temporal Alignment Connector
The results reveal that, compared to Libra-2, Libra-3 exhibits minimal changes in lexical scores, while clinical scores decline due to overfitting caused by the additional fine-tuning. Notably, the CheXpert (Smit et al., 2020a) (Macro-F1-[5/14]) scores exhibit the most influential reduction.
Despite this decline, ablation studies confirm that TAC’s performance improvements remain robust, unaffected by variations in training strategies. This resilience stems from TAC’s ability to capture and retain temporal image representations during the initial training phase, which are preserved through subsequent fine-tuning.
These findings underscore TAC’s reliability as a critical component for temporal information processing in RRG tasks. It ensures stability even under diverse training conditions.
H.5 Impact of Radiology-Specific Pre-trained Models on Libra